# bibliography.bib

### v1

@@ Abstract
Estimating treatment effects is crucial for personalized decision-making in medicine, but this task faces unique challenges in clinical practice. At training time, models for estimating treatment effects are typically trained on well-structured medical datasets that contain detailed patient information. However, at inference time, predictions are often made using textual descriptions (e.g., descriptions with self-reported symptoms), which are incomplete representations of the original patient information. In this work, we make three contributions. (1) We show that the discrepancy between the data available during training time and inference time can lead to biased estimates of treatment effects. We formalize this issue as an inference time text confounding problem, where confounders are fully observed during training time but only partially available through text at inference time. (2) To address this problem, we propose a novel framework for estimating treatment effects that explicitly accounts for inference time text confounding. Our framework leverages large language models together with a custom doubly robust learner to mitigate biases caused by the inference time text confounding. (3) Through a series of experiments, we demonstrate the effectiveness of our framework in real-world applications.
@@ 1 Introduction
Estimating the conditional average treatment effect (CATE) is crucial for personalized medicine [feuerriegel2024causal]. A growing number of machine learning methods have been developed for this purpose [[, e.g.,]]chipman2010bart,curth2021nonparametric,curth2021inductive,johansson2016learning,johansson2018learning,kunzel2019metalearners,kennedy2023towards,ma2024diffpo,ma2025foundation,nie2021quasi,shalit2017estimating,wager2018estimation. When estimating CATEs from observational data, it is necessary to account for confounders – i.e., variables that influence both treatment and outcome. If confounders are not properly adjusted for, the estimated treatment effects may not reflect the true causal effect and are thus biased [Pearl2009].
@@
Example (see Figure 1): Consider a health insurance provider developing a medical chatbot to assist with treatment recommendations [de2022design, xu2021chatbot]. The chatbot is trained on high-quality clinical datasets containing detailed patient histories, laboratory results, and physician assessments to ensure reliable predictions of treatment effects. However, at inference time, the chatbot relies primarily on free-text inputs from patients describing their symptoms, often without supporting diagnostic tests or clinical measurements. A similar challenge arises in emergency settings, such as during triage calls or ambulance dispatches, where only verbal symptom descriptions or brief text messages are available.
@@
To fill the above gap, we propose a novel framework for unbiased CATE estimation in the presence of inference time text confounding. We call our framework TCA (short for text confounding adjustment). Thereby, we aim to bridge the gap between advanced treatment effect estimation methods for text data and real-world medical constraints. Our TCA framework operates in three stages: 1 estimate nuisance functions and construct pseudo-outcome based on the true confounders; 2 generate surrogates of the text confounders by leveraging state-of-the-art large language models (LLMs); and 3 perform a doubly robust text-conditioned regression to obtain valid and unbiased CATE estimates. Crucially, we use LLMs not as causal reasoners – given their documented limitations in reliable inference [jin2023can] – but as a semantically rich text generator.
@@
Overall, our main contributions are the following:  [^1] Code is available at https://github.com/yccm/llm-tca. (1) We formalize the problem of inference time text confounding, which is common in real-world medical applications such as telemedicine, where full information about confounders is missing at inference time. (2) We develop a novel framework called TCA to adjust for inference time text confounding and provide unbiased treatment effect estimations. (3) We demonstrate the effectiveness and real-world applicability of our framework across extensive experiments.
@@ 2 Related work
CATE estimation: Numerous methods have been proposed for estimating the CATE [[, e.g.,]]johansson2016learning,johansson2018learning,kunzel2019metalearners,kennedy2023towards,nie2021quasi,shalit2017estimating,wager2018estimation. On the one hand, there are works that suggest specific model architectures to address the covariate shift between treated and control groups by learning balanced representations [johansson2016learning, shalit2017estimating]. However, such specialized model architectures are not designed to handle unstructured data such as text. On the one hand, meta-learners offer a more flexible approach for constructing CATE estimators that can incorporate arbitrary machine learning models (e.g., neural networks) [curth2021inductive, kunzel2019metalearners]. Common meta-learners include the S-learner [kunzel2019metalearners], T-learner [kunzel2019metalearners], and the DR-learner [kennedy2023towards]. Nevertheless, it is unclear how best to adapt meta-learners for complex settings such as those involving text data.
Of note, all of the above works operate in the standard CATE setting, where an identical set of confounders is observed at both training and inference time. However, if the confounding variables are different (e.g., if some confounders are only partially observed or hidden at inference time), standard methods will yield biased estimates that no longer reflect the true treatment effect [Pearl2009]. Motivated by this issue, we focus on a specific setting that is characterized by such discrepancy where confounders were fully observed at training time but are only partially available through unstructured text data at inference time.
Treatment effect estimation with text data: A growing body of research has investigated causal inference with text data [[, e.g.,]]chen2024proximal, daoud2022conceptualizing, deshpande2022deep, egami2022make, keith2020text, roberts2020adjusting, maiya2021causalnlp, mozer2020matching, gultchin2021operationalizing, keith2021text, veitch2020adapting, weld2022adjusting, zhang2020quantifying, zeng2022uncovering). Therein, text data can have various role: (a) text data can be treatment variable [[, e.g.,]]egami2022make, tan2014effect, maiya2021causalnlp, wood2018challenges, wang2019words, pryzant2021causal; (b) text data can be a mediator [[, e.g.,]]gultchin2021operationalizing, keith2021text, veitch2020adapting, zhang2020quantifying); and (c) text data can be an outcome variable [[, e.g.,]]gill2015judicial, sridhar2019estimating, koroleva2019measuring). In contrast to that, our setting considers text data as a confounder.
Several works have also explored (d) text data as a confounder [[, e.g.,]]chen2024proximal, daoud2022conceptualizing, keith2020text, roberts2020adjusting, mozer2020matching, weld2022adjusting, zeng2022uncovering). To adjust for text confounding, recent methods typically attempt to mitigate the underlying confounding bias by viewing text features as proxies for unobserved confounders [chen2024proximal, deshpande2022deep, roberts2020adjusting, veitch2020adapting].
@@
NLP for clinical decision support: Natural language processing (NLP) has become a useful tool for clinical decision support, such as by enabling to extract information from unstructured clinical texts [[, e.g.,]]chen2020trends, demner2009can,hossain2023natural,hossain2024natural,huang2019clinicalbert, zhu2020using, wu2020deep, salunkhe2021machine, hiremath2022enhancing). Early work focused on using NLP to process clinical notes and symptom descriptions for some basic analysis and classification tasks, while more recent approaches leverage deep learning for learning the representations, particularly transformer-based models like BERT and GPT [brown2020language, devlin2018bert, huang2019clinicalbert, dhawan2024end, lahat2024assessing]. In a similar vein, LLMs have shown promise in medical applications, such as clinical text summarization, diagnosis, and treatment recommendations [dhawan2024end, nazi2024large, spitzer2025effect, thirunavukarasu2023large, ullah2024challenges].
However, most NLP models – including LLMs – are designed for associative learning rather than causal reasoning [jin2023can, jin2023cladder]. Hence, LLM fails in addressing text confounding and would thus give biased estimates. This is unlike our TCA framework where we aim to give unbiased estimates.
LLMs for causal inference: Recent works have explored the capabilities of LLMs in causal reasoning across various causal tasks [[, e.g.,]]guo2024estimating, jin2023can, jin2023cladder, kiciman2023causal, zhang2023understanding, zevcevic2023causal. For example, [dhawan2024end] leverages LLMs to impute missing data for causal inference. However, recent works by [jin2023can, jin2023cladder] demonstrate that LLMs fail to reliably reason about causality, which raises concerns given the need for reliable inferences, particularly in medicine [Kern2025], where flawed causal reasoning of LLM can lead to harmful or even dangerous decisions.
@@
Differences to other settings: The main difference between our setting and other settings lies in the availability and role of confounders during training and inference time, see Table 1. (i) Comparison with the standard CATE estimation setting: Unlike the standard CATE settings [[, e.g.,]]kunzel2019metalearners,shalit2017estimating, where the same set of confounders is observed at both training and inference, our setting involves a true confounder $X$ that is observed during training but not inference (while some induced text confounder $T$ is unobserved during training but observed at inference). (ii) Comparison with the proxy-based setting: In the proxy-based setting [chen2024proximal, deshpande2022deep, roberts2020adjusting, veitch2020adapting], one views text $T$ as a noisy proxy of true confounder $X$, which implies that the true confounders $X$ should be assumed latent and are never directly observed, while the proxy $T$ should always be observed. By contrast, in our setting, $T$ is not a proxy but an induced variable from $X$, because $T$ encodes partial information of the confounders. $T$ is unobserved during training but fully observed during inference, whereas $X$ is observed during training but unobserved during inference, which is opposite to the proxy setting.
@@ 3 Problem setup
Clinical relevance: Our setting is highly relevant in medical practice. One application is telemedicine [haleem2021telemedicine] where CATE models are trained on well-curated medical datasets (e.g., trial data) containing comprehensive confounder information (where $X$ includes patient characteristics and detailed diagnostic features, different risk factors, e.g., age, gender, prior diseases). Hence, at training time, detailed information about confounders is available. In contrast, at inference time, detailed patient information is unavailable, simply because diagnostic facilities are not available in telemedicine. Hence, at inference time, the CATE model can only be applied to symptom descriptions from patients (where $T$ provides partial information about the patient characteristics and diagnostic features in $X$).
Potential outcomes framework: We build upon Neyman-Rubin potential outcomes framework [rubin2005causal]. Hence, $Y(a)\in\mathcal{Y}$ denotes the potential outcome for a treatment intervention $A=a$. We have two potential outcomes for each individual: $Y(1)$ if the treated (i.e., $A=1$), and $Y(0)$ if not treated (i.e., $A=0$). However, due to the fundamental problem of causal inference [holland1986statistics], only one of the potential outcomes is observed. Hence, $Y=AY(1)+\left(1-A\right)Y(0)$. The conditional average treatment effect (CATE) with respect to confounder $X$ is defined as $\tau(x)=\mathbb{E}[Y(1)-Y(0)\mid X=x]$, which is the expected treatment effect for an individual with covariate value $X=x$.
@@
Of note, the above assumptions are standard in causality [Pearl2009]. Consistency ensures that the observed outcome $Y$ aligns with the potential outcome $Y(a)$ under the assigned treatment $A=a$. Unconfoundedness implies there are no unobserved confounders, i.e., all factors influencing both the treatment and the outcome are captured in the observed covariates $X$. Overlap guarantees that the treatment assignment is non-deterministic, i.e., that every individual has a non-zero probability of receiving each treatment option.
In practice, it is natural to view the text as a noisy ‘measurement’ of the true confounder. We thus formalize the induced text confounder $T$ generated from $X$ by
@@ 4.1 Adjusting for inference time text confounding
How do naïve baselines work? A naïve baseline would attempt to directly estimate $\tau^{t}(t)$ from $\mathcal{D}_{T}$. We refer to this approach later as the text-based estimator (TBE) in the experiments. Assuming that $T$ would be observed during training, a model is trained to learn the response surface with respect to $\mu_{a}^{t}(t)$ that gives $\mathbb{E}[Y\mid A=a,T=t]$. Here, a naïve baseline directly estimates $\tau^{t}_{\text{naive }}(t)=\mathbb{E}[Y\mid A=1,T=t]-\mathbb{E}[Y\mid A=0,T=t]$.
Why do naïve baselines fail? A naïve method estimates $\tau^{t}_{\mathrm{naive}}(t)$ which is not equal to the true CATE $\tau^{t}(t)$. The induced text confounder $T$ generated from $X$ often does not capture all the necessary confounding information in $X$ for potential outcomes. As a result, there exists confounding information in the residual confounders $X\backslash T$ that affects both $A$ and $Y$, which are unobserved in $\mathcal{D}_{T}$. Due to the residual confounding, we have $\mathbb{E}[Y\mid A=a,T=t]\neq\mathbb{E}[Y(a)\mid T=t]$. Below, we state the pointwise confounding bias of the naïve baseline following [coston2020counterfactual].
@@ 4.3 Doubly-robust CATE estimation with text confounders
Here, we adapt a state-of-the-art CATE meta-learner for our framework. Specifically, we leverage a doubly-robust (DR) learner [kennedy2023towards] for estimating CATE $\tau^{t}(t)$ due to the favorable theoretical property of being double robust. This step has two sub-steps: (i) Nuisance functions estimation using the true confounder $X$ and pseudo-outcomes construction; and (ii) text-conditioned regression with pseudo-outcomes.
Estimating nuisance functions and pseudo-outcomes: First, we estimate the response surface $\mu_{a}^{x}(x)$ and the propensity score $\pi^{x}(x)$ using training data $\mathcal{D}_{X}$. Let $\hat{\eta}^{x}(x)=\left(\hat{\mu}_{0}^{x}(x),\hat{\mu}_{1}^{x}(x),\hat{\pi}^{x}(x)\right)$ denote the estimated nuisance functions, where $\hat{\mu}_{a}^{x}(x)$ is the estimated response surface for treatment $A=a$, $\hat{\pi}^{x}(x)$ is the estimated propensity score. Following [coston2020counterfactual], these estimates enable the construction of a doubly-robust pseudo-outcome for each observation sample via
@@
Text-conditioned regression: To estimate $\tau^{t}(t)$ using the auxiliary dataset $\tilde{\mathcal{D}}_{X}$ from Sec. 4.2, we map the text $\tilde{T}$ to the text embedding $\phi(\tilde{T})$ using a pretrained LLM. We then train a regression model $f_{\theta}:\mathcal{T}\rightarrow\mathbb{R}$ to predict the pseudo-outcome $\tilde{Y}$ from text embeddings $\phi(\tilde{T})$ by minimizing the loss

### v2

# LLM-Driven Treatment Effect Estimation Under Inference Time Text Confounding

@@ Abstract
Estimating treatment effects is crucial for personalized decision-making in medicine, but this task faces unique challenges in clinical practice. At training time, models for estimating treatment effects are typically trained on well-structured medical datasets that contain detailed patient information. However, at inference time, predictions are often made using textual descriptions (e.g., descriptions with self-reported symptoms), which are incomplete representations of the original patient information. In this work, we make three contributions. (1) We show that the discrepancy between the data available during training time and inference time can lead to biased estimates of treatment effects. We formalize this issue as an inference time text confounding problem, where confounders are fully observed during training time but only partially available through text at inference time. (2) To address this problem, we propose a novel framework for estimating treatment effects that explicitly accounts for inference time text confounding. Our framework leverages large language models (LLMs) together with a custom doubly robust learner to mitigate biases caused by the inference time text confounding. (3) Through a series of experiments, we demonstrate the effectiveness of our framework in real-world applications.
@@ 1 Introduction
Estimating the conditional average treatment effect (CATE) is crucial for personalized medicine [14]. A growing number of machine learning methods have been developed for this purpose [[, e.g.,]]chipman2010bart,curth2021nonparametric,curth2021inductive,johansson2016learning,johansson2018learning,kennedy2023towards,kunzel2019metalearners,ma2024diffpo,ma2025foundation,nie2021quasi,shalit2017estimating,wager2018estimation. When estimating CATEs from observational data, it is necessary to account for confounders – i.e., variables that influence both treatment and outcome. If confounders are not properly adjusted for, the estimated treatment effects may not reflect the true causal effect and are thus biased [48].
@@
Example (see Figure˜1): Consider a health insurance provider developing a medical chatbot to assist with treatment recommendations [55, 68]. The chatbot is trained on high-quality clinical datasets containing detailed patient histories, laboratory results, and physician assessments to ensure reliable predictions of treatment effects. However, at inference time, the chatbot relies primarily on free-text inputs from patients describing their symptoms, often without supporting diagnostic tests or clinical measurements. A similar challenge arises in emergency settings, such as during triage calls or ambulance dispatches, where only verbal symptom descriptions or brief text messages are available.
@@
To fill the above gap, we propose a novel framework for unbiased CATE estimation in the presence of inference time text confounding. We call our framework TCA (short for text confounding adjustment). Thereby, we aim to bridge the gap between advanced treatment effect estimation methods for text data and real-world medical constraints. Our TCA framework operates in three stages: 1 estimate nuisance functions and construct pseudo-outcome based on the true confounders; 2 generate surrogates of the text confounders by leveraging state-of-the-art large language models (LLMs); and 3 perform a doubly robust text-conditioned regression to obtain valid and unbiased CATE estimates. Crucially, we use LLMs not as causal reasoners – given their documented limitations in reliable inference [26] – but as a semantically rich text generator.
@@
Overall, our main contributions are the following:  [^1] Code is available at https://github.com/yccm/llm-tca (1) We formalize the problem of inference time text confounding, which is common in real-world medical applications such as telemedicine, where full information about confounders is missing at inference time. (2) We develop a novel framework called TCA to adjust for inference time text confounding and provide unbiased treatment effect estimations. (3) We demonstrate the effectiveness and real-world applicability of our framework across extensive experiments.
@@ 2 Related work
CATE estimation: Numerous methods have been proposed for estimating the CATE [[, e.g.,]]johansson2016learning,johansson2018learning,kennedy2023towards,kunzel2019metalearners,nie2021quasi,shalit2017estimating,wager2018estimation. On the one hand, there are works that suggest specific model architectures to address the covariate shift between treated and control groups by learning balanced representations [28, 54]. However, such specialized model architectures are not designed to handle unstructured data such as text. On the one hand, meta-learners offer a more flexible approach for constructing CATE estimators that can incorporate arbitrary machine learning models (e.g., neural networks) [7, 39]. Common meta-learners include the S-learner [39], T-learner [39], and the DR-learner [33]. Nevertheless, it is unclear how best to adapt meta-learners for complex settings such as those involving text data.
Of note, all of the above works operate in the standard CATE setting, where an identical set of confounders is observed at both training and inference time. However, if the confounding variables are different (e.g., if some confounders are only partially observed or hidden at inference time), standard methods will yield biased estimates that no longer reflect the true treatment effect [48]. Motivated by this issue, we focus on a specific setting that is characterized by such discrepancy where confounders were fully observed at training time but are only partially available through unstructured text data at inference time.
Treatment effect estimation with text data: A growing body of research has investigated causal inference with text data [[, e.g.,]]chen2024proximal, daoud2022conceptualizing, deshpande2022deep, egami2022make, gultchin2021operationalizing, keith2020text, keith2021text, roberts2020adjusting, maiya2021causalnlp, mozer2020matching, veitch2020adapting, weld2022adjusting, zhang2020quantifying, zeng2022uncovering). Therein, text data can have various roles: (a) text data can be treatment variable [[, e.g.,]]egami2022make, maiya2021causalnlp, pryzant2021causal, tan2014effect, wang2019words, wood2018challenges; (b) text data can be a mediator [[, e.g.,]]gultchin2021operationalizing, keith2021text, veitch2020adapting, zhang2020quantifying); and (c) text data can be an outcome variable [[, e.g.,]]gill2015judicial, koroleva2019measuring, sridhar2019estimating). In contrast to that, our setting considers text data as a confounder.
Several works have also explored (d) text data as a confounder [[, e.g.,]]chen2024proximal, daoud2022conceptualizing, keith2020text, mozer2020matching, roberts2020adjusting, weld2022adjusting,zeng2022uncovering). To adjust for text confounding, recent methods typically attempt to mitigate the underlying confounding bias by viewing text features as proxies for unobserved confounders [2, 10, 50, 61].
@@
NLP for clinical decision support: Natural language processing (NLP) has become a useful tool for clinical decision support, such as by enabling to extract information from unstructured clinical texts [[, e.g.,]]chen2020trends, demner2009can, huang2019clinicalbert, zhu2020using, wu2020deep, salunkhe2021machine, hiremath2022enhancing, hossain2023natural,hossain2024natural). Early work focused on using NLP to process clinical notes and symptom descriptions for some basic analysis and classification tasks, while more recent approaches leverage deep learning for learning the representations, particularly transformer-based models like BERT and GPT [1, 11, 12, 25, 40]. In a similar vein, LLMs have shown promise in medical applications, such as clinical text summarization, diagnosis, and treatment recommendations [12, 46, 56, 59, 60].
However, most NLP models – including LLMs – are designed for associative learning rather than causal reasoning [26, 27]. Hence, LLM fails in addressing text confounding and would thus give biased estimates. This is unlike our TCA framework where we aim to give unbiased estimates.
LLMs for causal inference: Recent works have explored the capabilities of LLMs in causal reasoning across various causal tasks [[, e.g.,]]guo2024estimating, jin2023can, jin2023cladder, kiciman2023causal, zevcevic2023causal, zhang2023understanding. For example, [12] leverages LLMs to impute missing data for causal inference. However, recent works by [26, 27] demonstrate that LLMs fail to reliably reason about causality, which raises concerns given the need for reliable inferences, particularly in medicine [35], where flawed causal reasoning of LLM can lead to harmful or even dangerous decisions.
@@
Differences to other settings: The main difference between our setting and other settings lies in the availability and role of confounders during training and inference time, see Table 1.
(i) Comparison with the standard CATE estimation setting: Unlike the standard CATE settings [[, e.g.,]]kunzel2019metalearners,shalit2017estimating, where the same set of confounders is observed at both training and inference, our setting involves a true confounder $X$ that is observed during training but not inference (while some induced text confounder $T$ is unobserved during training but observed at inference). (ii) Comparison with the proxy-based setting: In the proxy-based setting [2, 10, 50, 61], one views text $T$ as a noisy proxy of true confounder $X$, which implies that the true confounders $X$ should be assumed latent and are never directly observed, while the proxy $T$ should always be observed. By contrast, in our setting, $T$ is not a proxy but an induced variable from $X$, because $T$ encodes partial information of the confounders. $T$ is unobserved during training but fully observed during inference, whereas $X$ is observed during training but unobserved during inference, which is opposite to the proxy setting.
@@ 3 Problem setup
Clinical relevance: Our setting is highly relevant in medical practice. One application is telemedicine [19] where CATE models are trained on well-curated medical datasets (e.g., trial data) containing comprehensive confounder information (where $X$ includes patient characteristics and detailed diagnostic features, different risk factors, e.g., age, gender, prior diseases). Hence, at training time, detailed information about confounders is available. In contrast, at inference time, detailed patient information is unavailable, simply because diagnostic facilities are not available in telemedicine. Hence, at inference time, the CATE model can only be applied to symptom descriptions from patients (where $T$ provides partial information about the patient characteristics and diagnostic features in $X$).
@@
Potential outcomes framework: We build upon Neyman-Rubin potential outcomes framework [51]. Hence, $Y(a)\in\mathcal{Y}$ denotes the potential outcome for a treatment intervention $A=a$. We have two potential outcomes for each individual: $Y(1)$ if the treated (i.e., $A=1$), and $Y(0)$ if not treated (i.e., $A=0$). However, due to the fundamental problem of causal inference [22], only one of the potential outcomes is observed. Hence, $Y=AY(1)+\left(1-A\right)Y(0)$. The conditional average treatment effect (CATE) with respect to confounder $X$ is defined as $\tau(x)=\mathbb{E}[Y(1)-Y(0)\mid X=x]$, which is the expected treatment effect for an individual with covariate value $X=x$.
@@
Of note, the above assumptions are standard in causality [48]. Consistency ensures that the observed outcome $Y$ aligns with the potential outcome $Y(a)$ under the assigned treatment $A=a$. Unconfoundedness implies there are no unobserved confounders, i.e., all factors influencing both the treatment and the outcome are captured in the observed covariates $X$. Overlap guarantees that the treatment assignment is non-deterministic, i.e., that every individual has a non-zero probability of receiving each treatment option.
In practice, it is natural to view the text as a noisy ‘measurement’ of the true confounder. We thus formalize the induced text confounder $T$ generated from $X$.
@@ 4.1 Adjusting for inference time text confounding
How do naïve baselines work? A naïve baseline would attempt to directly estimate $\tau^{t}(t)$ from $\mathcal{D}_{T}$. Assuming that $T$ would be observed during training, a model is trained to learn the response surface with respect to $\mu_{a}^{t}(t)$ that gives $\mathbb{E}[Y\mid A=a,T=t]$. Here, a naïve baseline directly estimates $\tau^{t}_{\text{naive }}(t)=\mathbb{E}[Y\mid A=1,T=t]-\mathbb{E}[Y\mid A=0,T=t]$.
[…]
@@ 4.3 Doubly-robust CATE estimation with text confounders
Here, we adapt a state-of-the-art CATE meta-learner for our framework. Specifically, we leverage a doubly-robust (DR) learner [33] for estimating CATE $\tau^{t}(t)$ due to the favorable theoretical property of being double robust. This step has two sub-steps: (i) Nuisance functions estimation using the true confounder $X$ and pseudo-outcomes construction; and (ii) text-conditioned regression with pseudo-outcomes.
Estimating nuisance functions and pseudo-outcomes: First, we estimate the response surface $\mu_{a}^{x}(x)$ and the propensity score $\pi^{x}(x)$ using training data $\mathcal{D}_{X}$. Let $\hat{\eta}^{x}(x)=\left(\hat{\mu}_{0}^{x}(x),\hat{\mu}_{1}^{x}(x),\hat{\pi}^{x}(x)\right)$ denote the estimated nuisance functions, where $\hat{\mu}_{a}^{x}(x)$ is the estimated response surface for treatment $A=a$, $\hat{\pi}^{x}(x)$ is the estimated propensity score. Following [5], these estimates enable the construction of a doubly-robust pseudo-outcome for each observation sample via
@@
[…]